# TabDPT: Scaling Tabular Foundation Models on Real Data

### v2

@@ Abstract
Tabular data is one of the most ubiquitous sources of information worldwide, spanning a wide variety of domains. This inherent heterogeneity has slowed the development of Tabular Foundation Models (TFMs) capable of fast generalization to unseen datasets. In-Context Learning (ICL) has recently emerged as a promising solution for TFMs, enabling dynamic adaptation to new tasks without additional tuning. While many studies have attempted to re-purpose large language models for tabular ICL, they have had limited success, so recent works have focused on developing tabular-specific foundation models. In this work, we propose an approach to combine ICL-based retrieval with self supervised learning to train tabular foundation models. We also investigate the utility of real vs. synthetic data for model pre-training, and show that real data can contain useful signal not easily captured in synthetic training. Specifically, we show that incorporating real data during the pre-training phase can lead to significantly faster training and better downstream generalization to unseen data. Our resulting model, TabDPT, achieves top performance on both regression (CTR23) and classification (CC18) benchmarks. Importantly, we also demonstrate that with our pre-training procedure, scaling both model and data size leads to consistent performance improvements that follow power laws. This echoes scaling laws in LLMs and other foundation models, and suggests that Internet-scale TFMs can be achievable. We open-source our full pipeline: inference code including trained model weights can be found here, and the training code to reproduce experiments can be found here.
footnotetext: Equal Contribution
@@ 1 Introduction
We show that ICL retrieval combined with self-supervised learning (SSL) based on column masking [11, 25, 44] leads to a robust pre-training procedure for TFMs. We investigate the utility of real data, showing that applying this pre-training procedure to real data leads to faster convergence and improved downstream accuracy on unseen datasets compared to training exclusively on synthetic data. Since existing TFMs are predominately trained with synthetic data, our findings suggest that further investigation should be conducted into the benefits of curating real datasets for pre-training. Our pre-training process produces a TFM capable of both classification and regression with leading accuracy on new, unseen datasets without any further training or tuning. We name this new model the Tabular Discriminative Pre-trained Transformer, or TabDPT for short.
Comprehensive evaluations on the OpenML-CC18 [5] and OpenML-CTR23 [16] benchmarks confirm the effectiveness of TabDPT. It consistently matches or surpasses the performance of specialized models that undergo extensive per-dataset hyperparameter optimization at a fraction of the deployment time and cost. Furthermore, we show strong results in the few-shot regime, where, with minimal semi-supervised modifications, TabDPT outperforms specialized baselines on 10-shot classification tasks. Finally, we demonstrate that TabDPT scales predictably with both model size and quantity of real pre-training data (Figure 1), underscoring the viability of Internet-scale “foundation” models for tabular domains. We summarize our contributions as follows:
@@
    We show that applying our pre-training procedure to real data leads to faster convergence and better downstream accuracy than using purely synthetic data. We also demonstrate scaling with this procedure where more data and larger models continue to yield consistent gains akin to scaling laws in LLMs.
@@
    We release TabDPT – including open model weights, code, and pre-training datasets – to encourage further research and reproducibility. A lightweight inference interface is available at https://github.com/layer6ai-labs/TabDPT-inference, and the full training code at https://github.com/layer6ai-labs/TabDPT-training.
@@ 2 Related Work
Tabular Foundation Models   Although foundation models in other domains [11, 8, 12] have shown tremendous progress in recent years, foundation models for tabular data have lagged behind [55]. Several attempts have tried to bridge this gap. However, many of these methods require additional training when applied to a novel task [36, 56, 37], hindering widespread adoption in practice due to the high costs associated with fine-tuning. Meanwhile, ICL-based TFMs have started gaining traction. Among them, large language model (LLM) based approaches [17, 26, 53, 60, 61] initially appear to be a natural fit. However, LLMs cannot easily handle numerical content of tables since their tokenization is specifically designed for language data. As a result, LLM-based ICL methods suffer from high memory costs and low performance, as we discuss in Sections 3.1 and 4.3. Furthermore, natural language follows a sequential left-to-right structure, whereas tabular data is inherently order-invariant with respect to rows. Indeed, LLMs are sensitive to the order of examples in context [39]. On the other hand, tabular-specific ICL methods such as TabPFN [28] can naturally handle tabular data with numerical entries. However, they are completely reliant on synthetic data generators; ensuring that this mechanism captures the full diversity and nuances of real-world data is challenging, and making meaningful improvements to it is difficult. Notable concurrent work by Hollmann et al. [29] does not include open-sourced pre-training and synthetic generators, further complicating direct improvements to their model; another concurrent method requires a complex, three-stage procedure to learn from synthetic generators [47]. In this paper, we hypothesize that real tabular data contains much more information than heavily engineered synthetic tabular generators, thus allowing more straightforward improvements by scaling model and data size, which is supported by experiments in Section 4.5.
@@ 3.1 Tabular Transformer Encoder
We use a row-based transformer encoder similar to TabPFN [28], where each table row serves as a “token.” Specifically, for an input table with $N$ rows and $F$ features, we standardize its feature dimension to $F_{\max}$ via padding ($F<F_{\max}$) or dimensionality reduction ($F>F_{\max}$), then embed each row into a $d$-dimensional vector. Rows attend to each other through stacked transformer layers.
A key motivation behind row-based encoding is memory and compute efficiency. In typical cell- or text-based tokenizations [21, 56, 29], each cell in an $N\times F$ table must be split into multiple tokens (e.g., subwords), resulting in $\mathcal{O}(N\times F\times\langle N_{\text{tok}}\rangle)$ tokens, where $\langle N_{\text{tok}}\rangle$ is the average number of tokens per cell. Even modest-sized tables can inflate the input sequence well beyond typical transformer context limits. By contrast, encoding entire rows as tokens reduces the sequence length to $N$, allowing us to process many more rows with significantly lower memory overhead.
Finally, TabDPT uses a shared architecture for both regression and classification. This is realized through two separate MLP heads atop one single shared transformer: one head used for classification (supporting up to $C_{\max}$ classes) and another for regression. The shared backbone facilitates better parameter sharing across regression and classification tasks. Additional implementation details are provided in Appendix C.
@@ 3.2 Self-Supervised Learning on Tabular Data
Real-world tabular datasets are typically structured as $\mathcal{D}=\{X,y\}$, where $X\in\mathbb{R}^{N\times F}$ is the input table containing $N$ rows and $F$ features, and $y\in\mathbb{R}^{N\times 1}$ is the target (class index or regression value). In some instances $y$ can contain multiple targets but that is relatively rare: most publicly released datasets have a single target. As the number of high quality publicly available tabular datasets is also relatively low, training TFMs in a supervised fashion to predict $y$ quickly saturates and leads to overfitting. To circumvent this, current methods predominantly leverage synthetic data that is continuously generated throughout pre-training [28, 7, 29, 47]. However, this approach has its own set of challenges where priors that generate synthetic data need to be extensively engineered to approximate the distribution of highly heterogeneous real-world data. Moreover, very few synthetic data generators have been released, making it challenging to reproduce and advance efforts in this direction.
In this work we take a different approach and aim to maximize the value of real data in TFM pre-training. To this end, we leverage self-supervised learning (SSL) to extract maximal information from each table and regularize the model. Inspired by masked modeling in language [11] and vision [25], as well as promising results in tabular domain [44, 17], we randomly designate one column as the “target” to be predicted from the others. Concretely, we randomly pick a task to be either regression or classification, then pick a column $c$ with sufficient unique values as the target. We remove $c$ from the table and standardize its values for regression or bin them into classes for classification. The model then has to predict this auxiliary target $y=c$ from the resulting table $X\backslash c$.
We also shuffle and drop other columns, forcing the model to learn from varying feature combinations. Without these augmentations, the number of prediction tasks would grow only linearly with the number of features. In contrast, our approach scales task count combinatorially, compelling the model to capture richer inter-feature relationships. This provides a stronger training signal and serves simultaneously as a regularizer. Pseudo-code for the SSL procedure is provided in Appendix D.
@@ 3.3 Retrieval-Based Pre-Training
We propose to take this one step further and align training with inference by also leveraging retrieval during training batch construction. Formally, after obtaining $y=c$ and $X\backslash c$ through SSL, we sample a set of $K$ rows from $X\backslash c$ that are close to each other in the feature space. These rows are partitioned into two groups: “context” $\{X_{\text{ctx}},y_{\text{ctx}}\}$ and “query” $\{X_{\text{qy}},y_{\text{qy}}\}$. The context $\{X_{\text{ctx}},y_{\text{ctx}}\}$, together with the query features $X_{\text{qy}}$, is fed into the model to predict query targets $y_{\text{qy}}$. To form model inputs, we pass the context and query through the appropriate row embedding functions ($\phi_{x}$ or $\phi_{y}$), then sum the embeddings of $X_{\text{ctx}}$ and $y_{\text{ctx}}$, and concatenate with $X_{\text{qy}}$:
@@
Here, $\oplus$ is element-wise addition, and $\hat{y}_{\text{qy}}$ emerges from the classification or regression head depending on the target task. The notation [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] indicates the cross-attention split: query points attend to all context points, while context points attend to each other but not to queries—that is, only context serves as keys in the attention mechanism.
Using “similar” rows as context during training and limiting context size naturally aligns it with inference, improving model generalization. It also maintains efficient batch sizes, as context size no longer scales with table size, while still exposing the model to relevant context information. We consistently observe that training with this procedure speeds up convergence and leads to better downstream accuracy. An illustration of our model architecture is shown in Figure 2, with a full training step described in Algorithm 1.
@@ 4.1 Data
Training Data   Our training data was collected from OpenML [57] and consists of a wide range of public tabular datasets across diverse domains, all available under the CC-BY licence. To find appropriate datasets, we considered those specified in the Grinsztajn et al. [23], TabZilla [42], and AMLB [18] benchmarks, as well as additional datasets found individually. The full set of pre-training data contains 123 datasets, with a total of 32M rows and 2B cells (individual values within each table) from a diverse set of domains such as biology, finance, industrial applications, and medicine. The scale of this data is comparable to related work such as Tabula-8B [17], that fine-tuned the LLaMA 3-8B [22] language model for the tabular domain using real-world data. We conjecture that the diversity of domains present in our pre-training data can provide a salient signal and improve downstream generalization. Further details, including the complete list of training datasets and breakdown by size and domain are provided in Appendix B.
@@ 4.2 Baselines
We compare our method against leading baselines that are tuned for each dataset, including tree-based methods such as XGBoost [9] and CatBoost [46], and deep learning methods such as TabR [21] and MLP-PLR [20], as well as MLP. For XGBoost, CatBoost, and LightGBM, we use results reported in the TabZilla benchmark [42]. Some datasets are missing results, so we conduct hyperparameter optimization and train models following the TabZilla protocol using the code repository from [21] [^1] https://github.com/yandex-research/tabular-dl-tabr. For TabR, MLP-PLR, and MLP, we use the same code repository with the predefined search space and 30 search rounds for both CC18 and CTR23. We choose the best hyperparameters for each dataset fold individually based on the validation performance.
We also compare to ICL baselines including the LLM-based Tabula-8B [17], and tabular-specific foundation models TabPFN v2 [29], and TabPFN (kNN) [54], which retrieves neighbours of each query at inference time. We run all methods on at least two different splits of the data and report $95\%$ confidence intervals using bootstrapping [1]. For TabDPT, we use the 78M-parameter variant, with 16 transformer layers pre-trained for 600K steps. All training and inference is done on Nvidia A100 GPUs with 40 GB of memory. Further training details are provided in Appendix C
@@ 4.3 Results on CC18 and CTR23
AUC (rank) Accuracy (rank) Correlation (rank) $R^{2}$ (rank)
TFM TabDPT 2.826 ± [2.667, 2.986] 2.833 ± [2.701, 2.965] 2.829 ± [2.643, 3.014] 2.800 ± [2.629, 2.957]
TabPFN v2 [29] 3.479 ± [3.361, 3.597] 3.549 ± [3.368, 3.729] 3.057 ± [2.886, 3.229] 3.171 ± [2.957, 3.371]
TabPFN (kNN) [54] 5.403 ± [5.229, 5.583] 5.479 ± [5.292, 5.667] N/A N/A
Deep Learning TabR [21] 4.847 ± [4.583, 5.111] 4.000 ± [3.806, 4.194] 3.929 ± [3.714, 4.157] 4.014 ± [3.800, 4.243]
MLP-PLR [20] 4.792 ± [4.569, 5.014] 4.500 ± [4.278, 4.722] 3.914 ± [3.743, 4.086] 3.900 ± [3.743, 4.057]
MLP 7.340 ± [7.146, 7.528] 6.806 ± [6.632, 6.972] N/A N/A
Tree-Based XGBoost [9] 4.903 ± [4.694, 5.111] 4.438 ± [4.243, 4.625] 4.214 ± [4.000, 4.429] 4.114 ± [3.900, 4.329]
LightGBM [35] 5.194 ± [4.951, 5.438] 5.215 ± [5.014, 5.424] 4.629 ± [4.457, 4.800] 4.586 ± [4.414, 4.757]
CatBoost [46] 4.958 ± [4.743, 5.174] 4.910 ± [4.674, 5.146] 5.429 ± [5.271, 5.586] 5.414 ± [5.229, 5.600]
@@
Table 1: Main results comparing models on evaluation data. We report the average ranks across four metrics and their $95\%$ confidence intervals. The best algorithm for each metric is bolded. Tabula-8B [17] only reports results on a subset of datasets in CC18 so we conduct pairwise comparison against it on reported datasets in Figure 3(a).
@@
Our main results comparing models on the evaluation data are shown in Table 1. TabDPT shows the best overall performance across all models. It has an overlap in 95% confidence intervals with TabPFN v2 on CTR23, but significantly outperforms it on CC18. TabDPT also significantly outperforms both deep learning and tree-based methods that are trained for each dataset. These results indicate that real data can be effectively utilized with SSL to train robust TFMs with leading performance. We provide a breakdown of results in Section F.1, examining each algorithm’s performance under varying dataset sizes, numbers of features, categorical fraction, and percent missing. Results indicate that TabDPT is robust to dataset variations in all of these categories. For Table 1 we use TabDPT with $2,048$ context size and $8$ ensembles.
@@
Win-Rate Comparison   To get a more direct view of how various methods perform against each other, we compute pairwise win-rate statistics. We assign a “win” to the method if it achieves a higher accuracy score on CC18 or $R^{2}$ score on CTR23. This also allows us to compare against Tabula-8B on CC18 as it only reports results on 61 of 72 datasets. Figure 3(a) shows the win-rate matrix for all methods, painting a similar picture to Table 1: TabDPT performs best overall, followed by the strong tabular-based foundation model TabPFN v2. Tabula-8B with 32 shots – the leading LLM-based tabular foundation model – is not competitive with the other techniques across our benchmarks, indicating that current LLMs are not well adapted to the tabular domain and further techniques are needed. We extend the pairwise comparison between methods in Appendix E using both Elo [14] and Glicko2 [19] metrics, drawing similar conclusions to above.
Figure 3: (a) Pairwise win-rate comparison. A win is counted for the method that achieves the higher classification/regression accuracy/$R^{2}$ on a given dataset. (b) Inference runtime vs performance. TabDPT models are ordered by context size. Non-TFM baseline runtimes are the total of hyperparameter optimization and inference.

### v3

@@ Abstract
Tabular data is one of the most ubiquitous sources of information worldwide, spanning a wide variety of domains. This inherent heterogeneity has slowed the development of Tabular Foundation Models (TFMs) capable of fast generalization to unseen datasets. In-Context Learning (ICL) has recently emerged as a promising solution for TFMs, enabling dynamic adaptation to new tasks without additional tuning. While many studies have attempted to re-purpose large language models for tabular ICL, they have had limited success, so recent works have focused on developing tabular-specific foundation models. In this work, we propose an approach to combine ICL-based retrieval with self supervised learning to train tabular foundation models. We also investigate the utility of real vs. synthetic data for model pre-training, and show that real data can contain useful signal not easily captured in synthetic training. Specifically, we show that incorporating real data during the pre-training phase can lead to significantly faster training and better downstream generalization to unseen data. Our resulting model, TabDPT, achieves strong performance on both regression (CTR23) and classification (CC18) benchmarks. Importantly, we also demonstrate that with our pre-training procedure, scaling both model and data size leads to consistent performance improvements that follow power laws. This echoes scaling laws in LLMs and other foundation models, and suggests that large-scale TFMs can be achievable. We open-source our full pipeline: inference code including trained model weights can be found here, and the training code to reproduce experiments can be found here.
footnotetext: Equal Contribution; 1Correspondence to {valentin.t, anthony}@layer6.ai
@@ 1 Introduction
We show that ICL retrieval combined with self-supervised learning (SSL) based on column masking [11, 28, 48] leads to a robust pre-training procedure for TFMs. We investigate the utility of real data, showing that using it for this pre-training procedure leads to faster convergence and improved downstream accuracy on unseen datasets compared to training exclusively on competing open-source synthetic data generators. Since existing TFMs are predominately trained with synthetic data, our findings suggest that further investigation should be conducted into the benefits of curating real datasets for pre-training. Our pre-training process produces a TFM capable of both classification and regression with leading accuracy on new, unseen datasets without any further training or tuning. We name this new model the Tabular Discriminative Pre-trained Transformer, or TabDPT for short.
Comprehensive evaluations on the OpenML-CC18 [5] and OpenML-CTR23 [17] benchmarks confirm the effectiveness of TabDPT. It consistently matches or surpasses the performance of specialized models that undergo extensive per-dataset hyperparameter optimization at a fraction of the deployment time and cost. Furthermore, we show strong results in the few-shot regime, where, with minimal semi-supervised modifications, TabDPT outperforms specialized baselines on 10-shot classification tasks, highlighting its versatility. Finally, we demonstrate that TabDPT scales predictably with both model size and quantity of real pre-training data (Figure 1), underscoring the viability of large-scale foundation models for tabular domains. We summarize our contributions as follows:
@@
    We show that applying our pre-training procedure to real data leads to faster convergence and better downstream accuracy than using purely open-source synthetic data generators. We also demonstrate scaling with this procedure where more data and larger models continue to yield consistent gains akin to scaling laws in LLMs.
@@
    We release TabDPT – including open model weights, code, and pre-training datasets – to encourage further research and reproducibility. A lightweight inference interface is available here, and the full training code here.
Since its initial publication, TabDPT has also been compared externally against other tabular models – including other TFMs – on the popular TabArena [15] benchmark, available online here. The most recent results, as of January 14, 2026, demonstrate that TabDPT is the premier open source TFM:
  • TabDPT has the highest Elo on all tasks among fully open source TFMs, and is third overall on the leaderboard (closely behind RealMLP [33] and Real TabPFN-2.5 [26]; the former is not a TFM and the latter has closed source training).
  • TabDPT has the highest default performance on regression of any model, and is neck-and-neck with Real TabPFN-2.5 after tuning and ensembling.
@@ 2 Related Work
Tabular Foundation Models   Although foundation models in other domains [11, 8, 12] have shown tremendous progress in recent years, foundation models for tabular data have lagged behind [59]. Several attempts have tried to bridge this gap. However, many of these methods require additional training when applied to a novel task [39, 60, 41], hindering widespread adoption in practice due to the high costs associated with fine-tuning. Meanwhile, ICL-based TFMs have started gaining traction. Among them, large language model (LLM) based approaches [18, 29, 57, 64] initially appear to be a natural fit. However, LLMs cannot easily handle numerical content of tables since their tokenization is not tailor-made for numerical data. As a result, LLM-based ICL methods suffer from high memory costs and low performance, as we discuss in Sections˜3.1 and 4.3. Furthermore, natural language follows a sequential left-to-right structure, whereas tabular data is inherently order-invariant with respect to rows. Indeed, LLMs are sensitive to the order of examples in context [43].
On the other hand, tabular-specific ICL methods such as TabPFN [31] can naturally handle tabular data with numerical entries. However, they are completely reliant on synthetic data generators; ensuring that this mechanism captures the full diversity and nuances of real-world data is challenging, and making meaningful improvements to it is difficult. Notable concurrent work by Hollmann et al. [32] does not include open-sourced pre-training and synthetic generators, further complicating direct improvements to their model; another concurrent method requires a complex, three-stage procedure to learn from synthetic generators [51]. In this paper, we hypothesize that real tabular data contains much more information than existing heavily engineered synthetic tabular generators, thus allowing more straightforward improvements by scaling model and data size; this is supported by experiments in Section˜4.5.
@@ 3.1 Tabular Transformer Encoder
We use a row-based transformer encoder similar to TabPFN v1 [31], where each table row serves as a “token.” Specifically, for an input table with $N$ rows and $F$ features, we standardize its feature dimension to $F_{\max}$ via padding ($F<F_{\max}$) or dimensionality reduction ($F>F_{\max}$), then embed each row into a $d$-dimensional vector. Rows attend to each other through stacked transformer layers.
A key motivation behind row-based encoding is memory and compute efficiency. In cell- or text-based tokenizations [22, 60, 32], each cell in an $N\times F$ table must be split into multiple tokens (e.g., subwords), resulting in $\mathcal{O}(N\times F\times\langle N_{\text{tok}}\rangle)$ tokens, where $\langle N_{\text{tok}}\rangle$ is the average number of tokens per cell. Even modest-sized tables can inflate the input sequence well beyond typical transformer context limits. By contrast, encoding entire rows as tokens reduces the sequence length to $N$, allowing us to process many more rows with significantly lower memory overhead. It also affords invariance to row ordering (no positional encoding), which matches the perceived structure of tabular data. While cell-based tokenization additionally introduces an appealing invariance to column ordering, we can achieve it approximately by permuting columns at training time; other works [32, 51] have also found that strict column order invariance can introduce training instability.
Finally, TabDPT uses a shared architecture for both regression and classification, which has been concurrently explored in TabPFN v2 [32]. In our case, this is realized through two separate MLP heads atop one single shared transformer: one head used for classification (supporting up to $C_{\max}$ classes) and another for regression. The shared backbone facilitates better parameter sharing across regression and classification tasks. Additional implementation details are provided in Appendix˜C.
@@ 3.2 Self-Supervised Learning on Tabular Data
Real-world tabular datasets are typically structured as $\mathcal{D}=\{X,y\}$, where $X\in\mathbb{R}^{N\times F}$ is the input table containing $N$ rows and $F$ features, and $y\in\mathbb{R}^{N\times 1}$ is the target (class index or regression value). In some instances $y$ can contain multiple targets but that is relatively rare. As the number of high quality publicly available tabular datasets is also relatively low, training TFMs in a supervised fashion to predict $y$ quickly saturates and leads to overfitting. To circumvent this, current methods predominantly leverage synthetic data that is continuously generated throughout pre-training [31, 7, 32, 51]. However, this approach has its own set of challenges where priors that generate synthetic data need to be extensively engineered to approximate the distribution of highly heterogeneous real-world data.
[…]
We also shuffle and drop other columns, forcing the model to learn from varying feature combinations. Without these augmentations, the number of prediction tasks would grow only linearly with the number of features. In contrast, our approach scales task count combinatorially, compelling the model to capture richer inter-feature relationships. This provides a stronger training signal and serves simultaneously as a regularizer. Pseudo-code for the SSL procedure is provided in Appendix˜D.
@@ 3.3 Retrieval-Based Pre-Training
[…]
@@
Here, $\oplus$ is element-wise addition, and $\hat{y}_{\text{qy}}$ emerges from the classification or regression head depending on the target task. The notation $\cdot$$\cdot$] indicates the cross-attention split: query points attend to all context points, while context points attend to each other but not to queries – that is, only context serves as keys in the attention mechanism.
Using “similar” rows as context during training and limiting context size naturally aligns it with inference, improving model generalization. It also maintains efficient batch sizes, as context size no longer scales with table size, while still exposing the model to relevant context information. We consistently observe that training with this procedure speeds up convergence and leads to better downstream accuracy. An illustration of our model architecture is shown in Figure˜2, with a full training step described in Algorithm˜1.
@@ 4.1 Data
Training Data   Our training data was collected from OpenML [61] and consists of a wide range of public tabular datasets across diverse domains, all available under the CC-BY licence. To find appropriate datasets, we considered those specified in the Grinsztajn et al. [25], TabZilla [46], and AMLB [19] benchmarks, as well as additional datasets found individually. The full set of pre-training data contains 123 datasets, with a total of 32M rows and 2B cells (individual values within each table) from a diverse set of domains such as biology, finance, industrial applications, and medicine. The scale of this data is comparable to related work such as Tabula-8B [18] that fine-tuned the LLaMA 3-8B [24] language model for the tabular domain using real-world data. We conjecture that the diversity of domains present in our pre-training data can provide a salient signal and improve downstream generalization. Further details, including the complete list of training datasets and breakdown by size and domain are provided in Appendix˜B.
@@ 4.2 Baselines
We compare our method against leading baselines that are tuned for each dataset, including tree-based methods such as XGBoost [9], LightGBM [38], and CatBoost [50], and deep learning methods such as TabR [22], TabM [23] and MLP-PLR [21], as well as MLP. For XGBoost, CatBoost, and LightGBM, we use results reported in the TabZilla benchmark [46]. Some datasets are missing results, so we conduct hyperparameter optimization and train models following the TabZilla protocol using the code repository from [22]. [^1] https://github.com/yandex-research/tabular-dl-tabr For TabR, TabM, MLP-PLR, and MLP, we use the same code repository with the predefined search space and 30 search rounds for both CC18 and CTR23. We choose the best hyperparameters for each dataset fold individually based on the validation performance.
We also compare to ICL baselines including the LLM-based Tabula-8B [18], and tabular-specific foundation models TabPFN v2 [32], and TabPFN (kNN) [58], which retrieves neighbours of each query at inference time. Note that using retrieval with a cell-based method like TabPFN v2 adds prohibitive computational overhead. We run all methods on at least two different splits of the data and report $95\%$ confidence intervals using bootstrapping [1]. For TabDPT, we use the 78M-parameter variant, with 16 transformer layers pre-trained for 600K steps. All training and inference is done on Nvidia A100 GPUs with 40 GB of memory. Further training details are provided in Appendix˜C
@@ 4.3 Results on CC18 and CTR23
AUC Accuracy Correlation $R^{2}$
TabDPT 0.933 ± [0.929, 0.937] 0.884 ± [0.882, 0.887] 0.837 ± [0.826, 0.848] 0.742 ± [0.731, 0.754]
TabPFN v2 0.932 ± [0.928, 0.936] 0.872 ± [0.869, 0.875] 0.835 ± [0.825, 0.845] 0.740 ± [0.729, 0.751]
TabPFN (kNN) 0.918 ± [0.915, 0.921] 0.850 ± [0.847, 0.853] N/A N/A
TabPFN 0.898 ± [0.895, 0.901] 0.812 ± [0.810, 0.814] N/A N/A
TabDPT (no val.) 0.928 ± [0.924, 0.932] 0.874 ± [0.871, 0.877] 0.835 ± [0.826, 0.844] 0.739 ± [0.729, 0.749]
TabM 0.917 ± [0.913, 0.921] 0.878 ± [0.876, 0.881] 0.824 ± [0.815, 0.834] 0.732 ± [0.722, 0.742]
TabR 0.925 ± [0.922, 0.928] 0.874 ± [0.871, 0.877] 0.828 ± [0.814, 0.842] 0.714 ± [0.693, 0.735]
MLP-PLR 0.912 ± [0.905, 0.919] 0.869 ± [0.865, 0.872] 0.829 ± [0.821, 0.838] 0.716 ± [0.699, 0.732]
MLP 0.872 ± [0.867, 0.876] 0.809 ± [0.805, 0.813] N/A N/A
XGBoost 0.926 ± [0.922, 0.929] 0.869 ± [0.866, 0.872] 0.827 ± [0.818, 0.837] 0.711 ± [0.698, 0.725]
LightGBM 0.924 ± [0.920, 0.927] 0.862 ± [0.859, 0.866] 0.825 ± [0.816, 0.834] 0.713 ± [0.697, 0.729]
CatBoost 0.926 ± [0.923, 0.930] 0.864 ± [0.860, 0.867] 0.822 ± [0.808, 0.837] 0.703 ± [0.683, 0.723]
@@
Table 1: Main results comparing models on evaluation data. We report the average performance on four metrics and their $95\%$ confidence intervals. The best algorithm for each metric is bolded. The best model that does not append the validation set to the training set is underlined. Tabula-8B [18] only reports results on a subset of datasets in CC18 so we conduct pairwise comparison against it on reported datasets in Figure 3(a).
@@
Our main results comparing models on the evaluation data are shown in Table˜1. TabDPT shows the best overall performance across all models. It is competitive with TabPFN v2 across metrics, and significantly outperforms it on accuracy on CC18. TabDPT also significantly outperforms both deep learning and tree-based methods that are trained for each dataset. These results indicate that real data can be effectively utilized with SSL to train robust TFMs with leading performance. We provide a breakdown of results in Section˜F.1, examining each algorithm’s performance under varying dataset sizes, numbers of features, categorical fraction, and percent missing. Results indicate that TabDPT is robust to dataset variations in all of these categories. For Table˜1, we use TabDPT with 2,048 context size and 8 ensemble members, where randomness for a given ensemble member emerges by permuting the features and classes (if applicable).
Our foundation models require negligible hyperparameter tuning. Rather than reserving a validation set, we merge it with the training set. In contrast, several baselines perform hyperparameter tuning on a separate validation set (instead of cross-validation with a final refit on the combined train and validation set). To benchmark under the same constraint, we report a variant in which TabDPT receives only the training set, denoted by (no val.). This setting should be compared to classical methods; the default TabDPT setting is comparable to other foundation models that also access the validation set. Since TabDPT (no val.) never uses validation data, it is disadvantaged compared to methods that tune on validation data. Even so, TabDPT (no val.) remains competitive and is only outperformed by TabM [23] (using 30 hyperparameter tuning rounds) on CC18 accuracy.
@@
Win-Rate Comparison   To get a more direct view of how various methods perform against each other, we compute pairwise win-rate statistics. We assign a “win” to the method if it achieves a higher accuracy score on CC18 or $R^{2}$ score on CTR23. This also allows us to compare against Tabula-8B on CC18 as it only reports results on 61 of 72 datasets. Figure˜3(a) shows the win-rate matrix for all methods, painting a similar picture to Table˜1: TabDPT performs best overall, followed by the strong tabular-based foundation model TabPFN v2. Tabula-8B with 32 shots – the leading LLM-based tabular foundation model – is not competitive with the other techniques across our benchmarks, indicating that current LLMs are not well adapted to the tabular domain and further techniques are needed. We extend the pairwise comparison between methods in Appendix˜E using both Elo [14] and Glicko2 [20] metrics, drawing similar conclusions to above.
[…]